OpenStudy (anonymous):

On a trip Harry drove a steady speed for 3 hours. An accident slowed his speed by 30 mph for the last part of the trip. If the 190-mile trip took 4 hours, what was his speed during the first part of the trip?

OpenStudy (michele_laino):

how much time the first part of the trip has took?

OpenStudy (michele_laino):

@martaamador62

OpenStudy (anonymous):

They didn't mention it

OpenStudy (michele_laino):

I think it is 4-3 = 1 hour

OpenStudy (michele_laino):

do you agree?

OpenStudy (michele_laino):

oops.. I have made an error, the first time of the trip is three hours, whereas the second time of the trip is 4-3 = 1 hour

OpenStudy (shinalcantara):

remember the basic concept of speed.. v = d/t or you'll have: d = vt -------- Since it is known that the total distance is 190miles and that the speed during the 1st 3hours is unknown and given that the speed for the last hour is 30mph slower than the 1st speed then if we let x be the 1st speed we'll have: xt + (x-30)t = 190 since it took 3hrs for the 1st speed and 1 hr for the other you'll have: x(3) + (x-30)(1) = 190 3x + x - 30 = 190 3x + x = 190 + 30 4x = 220 x = 55mph

OpenStudy (anonymous):

oh i got it now, thank you :)

OpenStudy (shinalcantara):

yw :)